# On thermal transpiration and thermomolecular pressure difference

### v1

@@ 2.3. Proposition of $S_{\Omega}$ and $J$
Let $\Omega$ be a bounded convex domain with $C^{1}$ boundary, $\alpha\geq 0,\beta\in\mathbb{R}^{3},-3<\gamma\leq 1$ and $2\leq p\leq\infty$. Suppose that (1.6) holds. Then, the operator norm of $S^{\kappa}_{\Omega}:L^{p}_{\alpha,\beta}\to L^{p}_{\alpha,\beta+1}$ is $\mathcal{O}(\kappa)$.
@@
$\lVert J^{\kappa}\large(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+(P(f)+\phi_{2})\textbf{1}_{\partial\Omega^{-}_{2}})\rVert_{L^{2}_{0,\gamma/2}}\leq C\kappa\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert f\rVert_{L^{2}(\partial\Omega^{+}_{2})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{2})}\Big).$ (2.14)
@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
$\int_{\mathbb{S}^{2}}\langle v^{\prime}\rangle^{-\beta}\langle u^{\prime}\rangle^{-\beta}d\sigma\leq C_{\beta}\left(\frac{1}{\sqrt{1+|v|^{2}+|u|^{2}}}\right)^{-(\beta+1)},$ (2.15)
@@
$\int_{\mathbb{S}^{2}}\langle v^{\prime}\rangle^{-\beta}\langle u^{\prime}\rangle^{\beta}d\sigma=\int_{0}^{2\pi}\int_{0}^{\pi}\frac{\sin\theta}{\sqrt{A^{2}-B^{2}\cos^{2}\theta}}d\theta\phi.$
@@
The second-to-last line follows from the induction hypothesis, so (2.15) holds when $\beta\in\mathbb{N}$. For $\beta\notin\mathbb{N}$, Suppose that $\beta=n+d$ for some $n\in\mathbb{N}$ and $d\in(0,1)$. Then, we have
@@
Now, we provide an estimate of $Q_{+},Q_{-}$, which will be used in Chapter 5.
@@
Suppose that (1.6) holds. Then,
$\left|\int_{\mathbb{R}^{3}}|v|S^{\kappa}_{\Omega}\left(Q_{+}[f,f]\right)(x,v)dv\right|\leq C_{\gamma,\beta}\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}$ (2.16)
@@
$\left|\int_{\mathbb{R}^{3}}|v|S^{\kappa}_{\Omega}\left(Q_{-}[f,f]\right)(x,v)dv\right|\leq C_{\gamma,\beta}\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}$ (2.17)
@@
The proof of 2.17 follows directly from
$\lVert S^{\kappa}(f)\rVert_{L^{\infty}_{\alpha,\beta}(\Omega\times\mathbb{R}^{3})}\leq\frac{C}{|v|}\lVert f\rVert_{L^{\infty}_{\alpha,\beta}(\Omega\times\mathbb{R}^{3})}.$ (2.18)
So it is omitted here for simplicity of presentation. The proof of (2.16) proceeds as follows:
$\left|\int_{\mathbb{R}^{3}}|v|S^{\kappa}_{\Omega}\left(Q_{+}[f,f]\right)(x,v)dv\right|$ $\leq C\left|\int_{\mathbb{R}^{3}}\lVert Q_{+}[f,f]\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}dv\right|$
@@
$\gamma^{\varepsilon}\equiv\{(x,v)\in\gamma:|n(x)\cdot v|\geq\varepsilon,\ |v|\leq\tfrac{1}{\varepsilon}\}.$ (3.1)
@@
$\|f1_{\gamma^{\varepsilon}}\|_{L^{1}(\partial\Omega\times\mathbb{R}^{3})}\;\leq C_{\varepsilon,\Omega}\Big(\|f\|_{L^{1}(\Omega\times\mathbb{R}^{3})}+\|v\cdot\nabla_{x}f\|_{L^{1}(\Omega\times\mathbb{R}^{3})}\Big).$ (3.2)
Also, we recall Green’s identity. The proof can be found in [5].
@@
Assume that $f(x,v),\,g(x,v)\in L^{2}(\Omega\times\mathbb{R}^{3})$, $v\cdot\nabla_{x}f,\;v\cdot\nabla_{x}g\in L^{2}(\Omega\times\mathbb{R}^{3})$, and $f_{\gamma},\;g_{\gamma}\in L^{2}(\partial\Omega\times\mathbb{R}^{3})$. Then
$\iint_{\Omega\times\mathbb{R}^{3}}\bigl\{(v\cdot\nabla_{x}f)\,g+(v\cdot\nabla_{x}g)\,f\bigr\}\,dv\,dx=\int_{\gamma}fg\,d\gamma.$ (2.3)
We now state the main theorem of this chapter. The statement and strategy are similar to [14] and [13]. However, we use the Lemma 1.2 ([24]) to simplify the process. For convenience, we present the detail here.
@@
Let $\Omega\in\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary. Assume that $3<\gamma\leq 1$ and that (1.6) holds. Then, for any $\phi_{1}\in L^{2}_{0,-\gamma/2}(\Omega\times\mathbb{R}^{3})$, $\phi_{2}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{1})$, and $f_{in}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{2})$. There exists a unique solution $f\in L^{2}_{0,\gamma/2}$ to the following boundary value problem.
$\left\{\begin{aligned} v\cdot\nabla_{x}f+\frac{1}{\kappa}L(f)&=\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f&=P(f)+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (3.3)
@@
$\left\{\begin{aligned} \hat{\alpha}\nu(v)f_{k+1}+v\cdot\nabla_{x}f_{k+1}+\frac{1}{\kappa}L(f_{k+1})&=\phi_{1},&&\quad\text{in}\quad\in\Omega\times\mathbb{R}^{3},\\ f_{k+1}&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f_{k+1}&=\eta P(f_{k})+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (3.5)
where $0<\hat{\alpha},\eta<1$. By Green formula, we have
$\hat{\alpha}\lVert f_{k+1}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f_{k+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f_{k+1}),f_{k+1}\rangle$ (3.6)
@@
$J^{\kappa}\large(f_{in}\chi_{\partial\Omega^{-}_{1}}+(P(f_{k})+\phi_{2})\chi_{\partial\Omega^{-}_{2}})\in L^{2}_{0,\gamma/2}\qquad\text{if }\,f_{k}\in L^{2}_{\Omega^{+}_{2}}.$ (3.8)
We can then rewrite (3.5) in integral form as:
$(I-\frac{1}{\kappa}S^{\kappa}_{\Omega}K)f_{k+1}=J^{\kappa}\large(f_{in}\chi_{\partial\Omega^{-}_{1}}+(P(f_{k})+\phi_{2})\chi_{\partial\Omega^{-}_{2}})+S_{\Omega}\phi_{1}.$ (3.9)
@@
$\lVert\eta P(f_{k})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$ $=\eta^{2}\lVert P(f_{k})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+2\eta\langle P(f_{k+1}),\phi_{2}\rangle_{\partial^{-}_{2}}$ (3.10)
@@
Then, we combine the (3.6) and (3.10) to obtain
@@
$\quad+C_{\alpha}\left(\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}\right)+C_{\hat{\alpha},\eta}\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.$
@@
$\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $+(2\hat{\alpha})^{-1}\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ (3.12)
@@
$\hat{\alpha}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f^{\hat{\alpha},\eta}),f^{\hat{\alpha},\eta}\rangle$ (3.13)
@@
Applying (2.9) we have
$\lVert P_{L}(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\leq C\begin{pmatrix}[l]&\hat{\alpha}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}\\ &+\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.\end{pmatrix}$. (3.14)
Multiplying it with a constant $a>0$ and combining it with equation (3.13), we derive that
@@
$\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}$ $+\lVert(I-P)f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+C_{L}\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ (3.15)
@@
Hence, $f^{\hat{\alpha},\eta}$ is a bounded sequence in $L^{2}_{0,\gamma/2}$ for any fixed $\eta$. Observe that $f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{1}}$ satisfies
@@
$\lVert f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\quad+\lVert f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (3.16)
$\leq C_{\eta}\kappa\left(\hat{\alpha}_{1}\lVert f^{\eta,\hat{\alpha}_{1}}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\hat{\alpha}_{2}\lVert f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}\right)$
@@
This implies that $f^{\eta}=\lim_{\hat{\alpha}\to 0}f^{\hat{\alpha},\eta}$ exists in $L^{2}_{0,\gamma/2}$ for any $0<\eta<1$. Moreover, applying the same procedure as in 3, we obtain
$\lVert f^{\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\quad+\lVert f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+C_{L}\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}_{0,\gamma/2}}$ (3.17)
@@
$v\cdot\nabla_{x}(f^{\eta})^{2}=-2\frac{1}{\kappa}L(f^{\eta})f^{\eta}+2\phi_{1}f^{\eta}.$ (3.18)
Taking absolute values and integrating over $\Omega\times\mathbb{R}^{3}$ in (3.18), and using (3.17), we deduce that
@@
Using Lemma 3.1 , for any $\gamma^{\epsilon}$ in (3.1) away from $\partial\Omega_{0}$, we obtain
@@
$=\tfrac{1}{2}\lVert P(f^{\eta})\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}.$ (3.21)
@@
$\int_{|v|\geq 1/\varepsilon}m(v)\,|n(x)\cdot v|\,dv$ $\leq C\varepsilon.$ (3.22)
@@
$\lVert P(f^{\eta})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$ $=\lVert P(f^{\eta})\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (3.23)
$f^{\eta}$ $=P(f^{\eta})+(I-P)(f^{\eta})$ (3.24)
@@
$\lVert P(f^{\eta})\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}\leq 2\lVert f^{\eta}\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+2\lVert(I-P)(f^{\eta})\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}.$ (3.25)
Combining this estimate with 3.17, 3.20, 3 and 3.25, we conclude that

### v2

@@ 2.3. Proposition of $S_{\Omega}$ and $J$
Let $\Omega$ be a bounded convex domain with $C^{1}$ boundary, $\alpha\geq 0,\beta\in\mathbb{R}^{3},-3<\gamma\leq 1$ and $2\leq p\leq\infty$. Suppose that (8) holds. Then, the operator norm of $S^{\kappa}_{\Omega}:L^{p}_{\alpha,\beta}\to L^{p}_{\alpha,\beta+1}$ is $\mathcal{O}(\kappa)$.
We also have the following estimate.
@@
$\lVert J^{\kappa}\large(f_{in}\textbf{1}_{\partial\Omega^{-}_{1}}+(P(f)+\phi_{2})\textbf{1}_{\partial\Omega^{-}_{2}})\rVert_{L^{2}_{0,\gamma/2}}\leq C\kappa\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert f\rVert_{L^{2}(\partial\Omega^{+}_{2})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{2})}\Big).$ (45)
@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
$\int_{\mathbb{S}^{2}}\langle v^{\prime}\rangle^{-\beta}\langle u^{\prime}\rangle^{-\beta}d\sigma\leq C_{\beta}\left(\frac{1}{\sqrt{1+|v|^{2}+|u|^{2}}}\right)^{-(\beta+1)},$ (46)
@@
This statement can be found in [43]. We provide a proof below.
@@
$\int_{\mathbb{S}^{2}}\langle v^{\prime}\rangle^{-1}\langle u^{\prime}\rangle^{-1}d\sigma=\int_{0}^{2\pi}\int_{0}^{\pi}\frac{\sin\theta}{\sqrt{A^{2}-B^{2}\cos^{2}\theta}}d\theta\phi.$
@@
The second-to-last line follows from the induction hypothesis, so (46) holds when $\beta\in\mathbb{N}$. For $\beta\notin\mathbb{N}$, suppose that $\beta=n+d$ for some $n\in\mathbb{N}$ and $d\in(0,1)$. Then, we have
@@
Now, we provide estimates of $Q_{+},Q_{-}$, which will be used in section 5.
@@
Suppose that (8) holds. Then,
$\left|\int_{\mathbb{R}^{3}}|v|S^{\kappa}_{\Omega}\left(Q_{+}[f,f]\right)(x,v)dv\right|\leq C_{\gamma,\beta}\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}$ (47)
@@
$\left|\int_{\mathbb{R}^{3}}|v|S^{\kappa}_{\Omega}\left(Q_{-}[f,f]\right)(x,v)dv\right|\leq C_{\gamma,\beta}\lVert f\rVert^{2}_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}$ (48)
@@
The proof of (48) follows directly from
$\lVert S^{\kappa}(f)\rVert_{L^{\infty}_{\alpha,\beta}(\Omega\times\mathbb{R}^{3})}\leq\frac{C}{|v|}\lVert f\rVert_{L^{\infty}_{\alpha,\beta}(\Omega\times\mathbb{R}^{3})}.$ (49)
The proof of (47) proceeds as follows:
$\left|\int_{\mathbb{R}^{3}}|v|S^{\kappa}_{\Omega}\left(Q_{+}[f,f]\right)(x,v)dv\right|$ $\leq C\left|\int_{\mathbb{R}^{3}}\lVert Q_{+}[f,f]\rVert_{L^{\infty}_{0,\beta}(\Omega\times\mathbb{R}^{3})}dv\right|$
@@
$\gamma^{\varepsilon}\equiv\{(x,v)\in\partial\Omega\times\mathbb{R}^{3}:|n(x)\cdot v|\geq\varepsilon,\epsilon\leq\ |v|\leq\tfrac{1}{\varepsilon}\}.$ (50)
@@
$\|f1_{\gamma^{\varepsilon}}\|_{L^{1}(\partial\Omega\times\mathbb{R}^{3})}\;\leq C_{\varepsilon,\Omega}\Big(\|f\|_{L^{1}(\Omega\times\mathbb{R}^{3})}+\|v\cdot\nabla_{x}f\|_{L^{1}(\Omega\times\mathbb{R}^{3})}\Big).$ (51)
Also, we recall Green’s identity. Let $\gamma$ denote the trace operator.
@@
Assume that $f(x,v),\,g(x,v)\in L^{2}(\Omega\times\mathbb{R}^{3})$, $v\cdot\nabla_{x}f,\;v\cdot\nabla_{x}g\in L^{2}(\Omega\times\mathbb{R}^{3})$, and $\gamma f,\;\gamma g\in L^{2}(\partial\Omega\times\mathbb{R}^{3})$. Then
$\iint_{\Omega\times\mathbb{R}^{3}}\bigl\{(v\cdot\nabla_{x}f)\,g+(v\cdot\nabla_{x}g)\,f\bigr\}\,dv\,dx=\int_{\partial\Omega\times\mathbb{R}^{3}}\gamma f\gamma g\,dSdv.$ (2.3)
The proof can be found in [5]. We now state the main theorem of this section. The statement and strategy are similar to [13] and [14]. However, we need to modify the proof due to the difference between the mixed and diffuse boundary conditions. Moreover, we use Lemma 1.2 ([24]) to simplify the argument, and we introduce the damping term $\hat{\alpha}\nu f$ rather than $\hat{\alpha}f$ to prove Theorem 3.1 for both positive and negative $\gamma$. We present the details below.
@@
Let $\Omega\in\mathbb{R}^{3}$ be a bounded convex domain with $C^{2}$ boundary. Assume that $3<\gamma\leq 1$ and that (8) holds. Then, for any $\phi_{1}\in L^{2}_{0,-\gamma/2}(\Omega\times\mathbb{R}^{3})$, $\phi_{2}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{1})$, and $f_{in}\in L^{2}_{0,\gamma/2}(\partial\Omega^{-}_{2})$. There exists a unique solution $f\in L^{2}_{0,\gamma/2}$ to the following boundary value problem.
$\left\{\begin{aligned} v\cdot\nabla_{x}f+\frac{1}{\kappa}L(f)&=\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f&=P(f)+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (52)
@@
$\left\{\begin{aligned} \hat{\alpha}\nu(v)f_{k+1}+v\cdot\nabla_{x}f_{k+1}+\frac{1}{\kappa}L(f_{k+1})&=\phi_{1},&&\quad\text{in}\quad\in\Omega\times\mathbb{R}^{3},\\ f_{k+1}&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f_{k+1}&=\eta P(f_{k})+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2},\end{aligned}\right$. (54)
where $0<\hat{\alpha},\eta<1$. Note that (54) and (52) are slightly different. The boundary value problem (54) includes a damping term $\hat{\alpha}\nu(v)f_{k+1}$ and we insert $\eta$ in front of $P(f)$. We first construct a solution of (54) in $L^{2}_{0,\gamma/2}$ for arbitrary $\hat{\alpha}$ and $\eta$. Let $f^{\hat{\alpha},\eta}$ denote this solution. We then show that, for fixed $\eta$, the family $\{f^{\hat{\alpha},\eta}\}$ converges in $L^{2}_{0,\gamma/2}$. Finally, we remove $\eta$ by proving that $\{f^{\eta}\}$ is a convergent sequence in $L^{2}_{0,\gamma/2}$,where $f^{\eta}=\lim_{\hat{\alpha}\to 0}f^{\hat{\alpha},\eta}$. To construct the solution of (54), by Green formula, we have
$\hat{\alpha}\lVert f_{k+1}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f_{k+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f_{k+1}),f_{k+1}\rangle$ (55)
@@
$J^{\kappa}\large(f_{in}\chi_{\partial\Omega^{-}_{1}}+(P(f_{k})+\phi_{2})\chi_{\partial\Omega^{-}_{2}})\in L^{2}_{0,\gamma/2}\qquad\text{if }\,f_{k}\in L^{2}_{\Omega^{+}_{2}}.$ (57)
We can then rewrite (54) in integral form as:
$(I-\frac{1}{\kappa}S^{\kappa}_{\Omega}K)f_{k+1}=J^{\kappa}\large(f_{in}\chi_{\partial\Omega^{-}_{1}}+(P(f_{k})+\phi_{2})\chi_{\partial\Omega^{-}_{2}})+S_{\Omega}\phi_{1}.$ (58)
@@
$\lVert\eta P(f_{k})+\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$ $=\eta^{2}\lVert P(f_{k})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}+2\eta\langle P(f_{k+1}),\phi_{2}\rangle_{\partial^{-}_{2}}$ (59)
@@
Then, we combine the (55) and (59) to obtain
@@
$\quad+C_{\hat{\alpha},\eta}\left(\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}\right)+C_{\hat{\alpha},\eta}\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.$
@@
$\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $+(2\hat{\alpha})^{-1}\lVert f_{n+1}-f_{m+1}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ (61)
@@
$\hat{\alpha}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\frac{1}{2}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+})}$ $+\frac{1}{\kappa}\langle L(f^{\hat{\alpha},\eta}),f^{\hat{\alpha},\eta}\rangle$ (62)
@@
Applying (35) we have
$\lVert P_{L}(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\leq C\begin{pmatrix}[l]&\hat{\alpha}\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}\\ &+\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}.\end{pmatrix}$. (63)
Multiplying it with a constant $a>0$ and combining it with equation (62), we derive that
@@
$\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}$ $+\lVert(I-P)f^{\hat{\alpha},\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+C_{L}\lVert(I-P_{L})(f^{\hat{\alpha},\eta})\rVert^{2}_{L^{2}_{0,\gamma/2}}$ (64)
@@
Hence, $f^{\hat{\alpha},\eta}$ is a bounded sequence in $L^{2}_{0,\gamma/2}$ for any fixed $\eta$. Observe that $f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}$ satisfies
@@
$\lVert f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\quad+\lVert f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)\Big(f^{\eta,\hat{\alpha}_{1}}-f^{\eta,\hat{\alpha}_{2}}\Big)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (65)
$\leq C_{\eta}\cdot\max\{1,\kappa\}\left(\hat{\alpha}_{1}\lVert f^{\eta,\hat{\alpha}_{1}}\rVert^{2}_{L^{2}_{0,\gamma/2}}+\hat{\alpha}_{2}\lVert f^{\eta,\hat{\alpha}_{2}}\rVert^{2}_{L^{2}_{0,\gamma/2}}\right)$
@@
This implies that $f^{\eta}=\lim_{\hat{\alpha}\to 0}f^{\hat{\alpha},\eta}$ exists in $L^{2}_{0,\gamma/2}$ for any $0<\eta<1$. Moreover, applying the same procedure as in (3), we obtain
$\lVert f^{\eta}\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\quad+\lVert f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert(I-P)f^{\eta}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+C_{L}\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}_{0,\gamma/2}}$ (66)
@@
$v\cdot\nabla_{x}(f^{\eta})^{2}=-\frac{2}{\kappa}L(f^{\eta})f^{\eta}+2\phi_{1}f^{\eta}.$ (67)
Taking absolute values and integrating over $\Omega\times\mathbb{R}^{3}$ in (67), and using (66), we deduce that
@@
Using Lemma 3.1 , for any $\gamma^{\epsilon}$ in (50) away from $\partial\Omega_{0}$, we obtain
@@
$=\tfrac{1}{2}\lVert P(f^{\eta})\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}.$ (70)
@@
$\int_{|v|\geq 1/\varepsilon}m(v)\,|n(x)\cdot v|\,dv$ $\leq C\varepsilon.$ (71)
@@
$\lVert P(f^{\eta})\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}$ $=\lVert P(f^{\eta})\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}$ (72)
$f^{\eta}$ $=P(f^{\eta})+(I-P)(f^{\eta})$ (73)
@@
$\lVert P(f^{\eta})\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}\leq 2\lVert f^{\eta}\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+2\lVert(I-P)(f^{\eta})\textbf{1}_{\gamma^{\epsilon}}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}.$ (74)
Combining this estimate with (66), (69), (3) and (74), we conclude that